# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ 5.1. Preliminaries
Corollaries 5.5 and 5.6 were proved in Győry (1976) with an explicit upper bound for the sizes of the solutions, not only for equations (5.6) and (5.7) but also for index form equations related to indices with respect to arbitrary orders $\mathcal{O}$ of $K$; see also Győry (2000) and Evertse and Győry (2017).
@@ 6. Algorithmic resolution of index form equations, application to (multiply) monogenic number fields
The exponential bound (5.8) for the solutions of (5.7) is too large for practical use. In the 1990’s, there were new breakthroughs, leading to the complete resolution of certain index form equations. In fact, efficient methods were elaborated for solving equation (5.7) when $|D_{K}|$ is not too large, and the degree $n$ of $K$ is $\leq 6$. Further, (5.7) was solved for many special higher degree number fields $K$ up to about degree $15$ and for some relative extensions of degree $\leq 4$.
@@
In some cases, the solution of index form equations such as (5.7) can be reduced to one or more Thue equations, i.e., equations of the form $F(x,y)=m$ in $x,y\in\mathbb{Z}$, where $F$ is a binary form with integer coefficients, and $m$ is a non-zero integer. Such an equation is called cubic if $F$ has degree $3$ and quartic if $F$ has degree $4$. For such equations there are efficient algorithms for solving the equations, provided that $H(F)$ and $|m|$ are not too large; see e.g. the books Smart (1998) or Gaál (2019).
For $\mathbf{n=3}$, Gaál and Schulte (1989) reduced (5.7) to a cubic Thue equation. Then, using an efficient algorithm for solving cubic Thue equations, they determined all the solutions of cubic index form equations (5.7), provided that $|D_{K}|$ is not too large.
For $\mathbf{n=4}$, Gaál, Pethő and Pohst (1993, 1996) reduced equation (5.7) to one cubic and some quartic Thue equations. Then, by means of efficient algorithms for solving such Thue equations, they computed all the solutions of equation (5.7) for quartic number fields with not too large discriminant. They obtained several very interesting related results as well.
6.2. Refined version of the general approach combined with reduction and enumeration algorithms
@@
For $\mathbf{n\geq 5}$, the approach via Thue equations does not work. In general, for $\mathbf{n=5}$ and $\mathbf{6}$ a refined version of the general approach involving unit equations is needed. Since by (5.5), (5.2a) and (5.1) we have
(5.7)$\Leftrightarrow D_{K/\mathbb{Q}}(\alpha)=D_{K}\Leftrightarrow D(f_{\alpha})=D_{K}\ \text{in\ }\alpha\in\mathcal{O}_{K}$
where $f_{\alpha}\in\mathbb{Z}[X]$ is the minimal polynomial of $\alpha$, in case of concrete equations (5.7) the basic idea of the proof of Theorem 4.2 for irreducible $f_{\alpha}$’s must be combined with some reduction and enumeration algorithms.
The refined version of the general method in the irreducible case consists of the following steps:
  1. (1)
    Reduction to unit equations but in considerably smaller subfields of the normal closure $G$ of $K$. Then the number $r$ of unknown exponents $a_{ijk,l}(1\leq l\leq r)$ in the unit equation (4.5) with $\varepsilon_{ijk}=\zeta_{ijk}\rho_{1}^{a_{ijk,1}}\cdots\rho_{r}^{a_{ijk,r}}$ is much smaller, at most $n(n-1)/2-1$ instead of $r\leq n!-1$; cf. Győry (1998, 2000). Then, in concrete cases one can bound the exponents $|a_{ijk,l}|$ by Baker’s method.
  2. (2)
    The bounds in concrete cases are still too large. Hence a reduction algorithm is needed, reducing the Baker’s bound for $|a_{ijk,l}|$ in several steps if necessary by refined versions of the $L^{3}$-algorithm; cf. de Weger (1989), Wildanger (1997) and Gaál and Pohst (1996).
  3. (3)
@@
Combining the refined version of the general method with reduction and enumeration algorithms, for $n=5,6$ and for not too large $|D_{K}|$, Gaál and Győry (1999), resp. Bilu, Gaál and Győry (2004) gave algorithms for determining all power integral bases and hence checking the monogenicity and the multiplicity of the monogenicity of $K$.
@@
In the examples below, the authors resolved concrete index form equations of the form (5.7) for $n=3,4,5,6$. The involved number fields $K$ of degree $n$ are given by irreducible monic polynomials $f(X)\in\mathbb{Z}[X]$, a zero of which generates the corresponding $K$ over $\mathbb{Q}$. In each case all power integral bases in $K$, and therefore the multiplicity of the monogenicity of $K$, denoted by $mm(K)$, are computed by the method outlined above. For the lists of the power integral bases, we refer to the original papers and to Evertse and Győry (2017) and Gaál (2019).
@@
  • $\mathbf{n=6}$, $f(X)=X^{6}-5X^{5}+2X^{4}+18X^{3}-11X^{2}-19X+1$, $mm(K)=45$ (Bilu, Gaál, and Győry, 2004);
We note that from the point of view of computation, the above examples belong to the most difficult cases, $K$ being in each case totally real with Galois group $S_{n}$. In these cases the number of exponents in the unit equations involved is the largest possible.
For $n\geq 7$, the above mentioned algorithms do not work in general. Then the number of fundamental units, $\rho_{1},\ldots,\rho_{r}$ involved can be too large to use the enumeration algorithm.
@@
For $\mathbf{n=7}$, give a practical algorithm for solving equation (5.7) in case of any number field $K$ of degree $7$ with not too large discriminant.

### v2

@@ 5.3. Consequences for monogenic number fields and orders
5.5. A consequence of Theorem 4.4 for algebraic numbers of given discriminant
a
Theorem 4.4 can be applied to algebraic numbers that are not necessarily algebraic integers. Given an algebraic number $\alpha$, we denote by $f_{\alpha}$ its primitive minimal polynomial, i.e.,
(5.11) $f_{\alpha}=a_{0}X^{n}+\cdots+a_{n}=a_{0}(X-\alpha^{(1)})\cdots(X-\alpha^{(n)})\in\mathbb{Z}[X]$
where $a_{0}>0$, ${\rm gcd}(a_{0},\ldots,a_{n})=1$ and $\alpha^{(1)}=\alpha,,\ldots,\alpha^{(n)}$ are the distinct conjugates of $\alpha$. We recall that the height and discriminant of $\alpha$ are defined by those of $f_{\alpha}$, i.e.,
$H(\alpha):=H(f_{\alpha}),\ \ D(\alpha):=D(f_{\alpha}).$
Two algebraic numbers $\alpha,\beta$ are called $GL_{2}(\mathbb{Z})\textit{-equivalent}$ if
$\beta=\frac{a\alpha+b}{c\alpha+d}\text{\,\, with \,}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in GL_{2}(\mathbb{Z}).$
One easily verifies that if $\alpha,\beta$ are $GL_{2}(\mathbb{Z})$-equivalent then so are $f_{\alpha}$, $f_{\beta}$ while conversely, if $f_{\alpha},f_{\beta}$ are $GL_{2}(\mathbb{Z})$-equivalent, then $\alpha$ is $GL_{2}(\mathbb{Z})$-equivalent to a conjugate of $\beta$.
Consequently, if $\alpha$, $\beta$ are $GL_{2}(\mathbb{Z})$-equivalent, then $D(\alpha)=D(\beta)$. Now Theorem 4.4 implies at once:
Theorem 5.10 (Evertse and Győry, 1991a).
Every algebraic number $\alpha$ of degree $n\geq 2$ and discriminant $D\neq 0$ is $GL_{2}(\mathbb{Z})$-equivalent to an algebraic number $\beta$ with
$H(\beta)\leq c_{3}(n,|D|),$
where $c_{3}$ denotes the same effectively computable positive number as in Theorem 4.4.
Further, by Théorème 1 of Győry (1974) we have
$n\leq 2\log|D|/\log 3.$
5.6. Rationally monogenic orders
a
Monogenic orders $\mathbb{Z}[\alpha]$, where $\alpha$ is an algebraic integer, can be generalized to so-called rationally monogenic orders $\mathbb{Z}_{\alpha}$, where $\alpha$ is not necessarily integral. We will formulate an analogue of Corollary 5.3 for rationally monogenic orders. While in the results for monogenic orders, $\mathbb{Z}$-equivalence of algebraic integers plays an important role, for rationally monogenic orders we have to deal with $GL_{2}(\mathbb{Z})$-equivalence of algebraic numbers. Before we define rationally monogenic orders, we briefly go into some history and introduce the necessary terminology.
Let $\alpha$ be a non-zero, not necessarily integral algebraic number of degree $n\geq 3$, and $f_{\alpha}$ its primitive minimal polynomial, given by (5.11). Define $\mathbb{Z}_{\alpha}$ to be the $\mathbb{Z}$-module with basis
This $\mathbb{Z}$-module was introduced by Birch and Merriman (1972), who observed that it is contained in the ring of integers of $\mathbb{Q}(\alpha)$, and that for its discriminant we have
(5.12) $D(\mathbb{Z}_{\alpha})=D(f_{\alpha})=D(\alpha).$
Nakagawa (1989) showed that $\mathbb{Z}_{\alpha}$ is in fact an order of the field $\mathbb{Q}(\alpha)$, i.e., closed under multiplication. More precisely, he showed that
(5.13) $\omega_{i}\omega_{j}=-\sum_{\max(i+j-n,1)\leq k\leq i}a_{i+j-k}\omega_{k}+\sum_{j<k\leq\min(i+j,n)}a_{i+j-k}\omega_{k}$
for $i,j=1,\ldots,n-1$, where $\omega_{n}:=-a_{n}$. This order was further studied by Simon (2001, 2003) and Del Corso, Dvornicich and Simon (2005). They showed that
(5.14) $\mathbb{Z}_{\alpha}=\mathbb{Z}[\alpha]\cap\mathbb{Z}[\alpha^{-1}].$
As was very likely known at the time, another description of $\mathbb{Z}_{\alpha}$ is as follows. Let $\mathcal{M}_{\alpha}$ be the $\mathbb{Z}$-module generated by $1,\alpha,\ldots,\alpha^{n-1}$. Then $\mathbb{Z}_{\alpha}$ is the ring of coefficients of $\mathcal{M}_{\alpha}$ (see Borevich and Shafarevich (1967), Section 2.2), i.e.,
(5.15) $\mathbb{Z}_{\alpha}=\left\{\xi\in\mathbb{Q}(\alpha):\ \xi\mathcal{M}_{\alpha}\subseteq\mathcal{M}_{\alpha}\right\}.$
We have
(5.16) $\mathbb{Z}_{\alpha}=\mathbb{Z}[\alpha]\ \ \text{if $\alpha$ is an algebraic integer.}$
Indeed, if $\alpha$ is an algebraic integer of degree $n$, the powers $\alpha^{i}$ ($i\geq n$) belong to $\mathcal{M}_{\alpha}$, and thus, $\mathbb{Z}_{\alpha}=\mathcal{M}_{\alpha}=\mathbb{Z}[\alpha]$. Further, for any two non-zero algebraic numbers $\alpha,\beta$ we have
[…]
Indeed, let $\beta=\mbox{\large{$\textstyle{\frac{a\alpha+b}{c\alpha+d}}$}}$ for some matrix $\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)\in GL_{2}(\mathbb{Z})$. Then $\mathcal{M}_{\beta}=(c\alpha+d)^{1-n}\mathcal{M}_{\alpha}$ where $n=\deg\alpha$, and thus, $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$.
We call an order $\mathcal{O}$ of a number field $K$ rationally monogenic if there is $\alpha$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$. From (5.16) it follows that monogenic orders are rationally monogenic. Below we explain that rationally monogenic orders are in fact special cases of invariant orders of polynomials. In particular, $\mathbb{Z}_{\alpha}$ is the invariant order of $f_{\alpha}$.
Recall that the index of an algebraic integer was defined in (5.4). Following Simon (2001), we generalize this to not necessarily integral algebraic numbers as follows. Given a non-zero algebraic number $\alpha$, we define the index of $\alpha$ by
$I(\alpha):=[\mathcal{O}_{K}:\mathbb{Z}_{\alpha}],$
where $K=\mathbb{Q}(\alpha)$. In fact, this is the index of $f_{\alpha}$ as it was introduced by Simon. From (5.1) and (5.12) we deduce, analogously to the first identity in (5.5),
(5.18) $D(\alpha)=I(\alpha)^{2}D_{K}.$
For more results and properties of this index, we refer to Simon (2001).
There is a connection between rationally monogenic orders and Hermite equivalence classes of polynomials, which we explain here without proof. For a non-zero algebraic number $\alpha$, let $\mathcal{I}_{\alpha}$ be the fractional ideal of $\mathbb{Z}_{\alpha}$ generated by $1$ and $\alpha$. This is known to be invertible, see Simon (2003). It is called also the invariant ideal of $f_{\alpha}$.
Theorem 5.11 (BEGyRS, 2023).
Let $f,g\in\mathbb{Z}[X]$ be two primitive, irreducible polynomials. Then the following three assertions are equivalent:
  • (i)
    $f$ and $g$ are Hermite equivalent;
  • (ii)
    $f$ has a root $\alpha$ and $g$ a root $\beta$ such that $\mathcal{M}_{\beta}=\lambda\mathcal{M}_{\alpha}$ for some non-zero $\lambda\in\mathbb{Q}(\alpha)$;
  • (iii)
In the particular case that $f$ and $g$ are monic, we have $\alpha\in\mathbb{Z}[\alpha]=\mathbb{Z}_{\alpha}$ and $\mathcal{I}_{\alpha}=\mathbb{Z}_{\alpha}$ and likewise for $g$ and $\beta$. This leads to the following corollary.
Corollary 5.12 (BEGyRS, 2023).
Let $f,g\in\mathbb{Z}[X]$ be two monic, irreducible polynomials. Then $f$ and $g$ are Hermite equivalent if and only if $f$ has a root $\alpha$ and $g$ a root $\beta$ such that $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$.
In BEGyRS (2023) an example of a quartic algebraic number field $K$ was given such that $\mathcal{O}_{K}=\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$ for certain $\alpha,\beta\in K$, but $f_{\alpha}$, $f_{\beta}$ lie in different Hermite equivalence classes. So far, we haven’t been able to find similar examples for algebraic number fields of degree $\geq 5$.
An order $\mathcal{O}$ of a number field $K$ is called primitive if there are no integer $a>1$ and order $\mathcal{O}^{\prime}$ such that $\mathcal{O}=\mathbb{Z}+a\mathcal{O}^{\prime}$. It is not difficult to show that a rationally monogenic order is primitive. It follows from work of Delone and Faddeev (1940) that every primitive order of a cubic number field is rationally monogenic. Simon (2001) gave various examples of number fields of degree $\geq 4$ that are not rationally monogenic, i.e., whose rings of integers are not rationally monogenic.
In Evertse (2023) the following was shown:
Theorem 5.13.
Every number field $K$ of degree $\geq 3$ has infinitely many orders that are rationally monogenic but not monogenic.
We finally arrive at the main result of this subsection, which follows directly from Theorem 5.10 and (5.12):
Theorem 5.14.
Let $\mathcal{O}$ be an order of a number field $K$, and denote by $D(\mathcal{O})$ its discriminant. Then every $\alpha$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$ is $GL_{2}(\mathbb{Z})$-equivalent to some $\beta\in K$ of height $H(\beta)\leq c_{3}(n,|D(\mathcal{O})|)$, where $c_{3}$ denotes the same effectively computable positive number as in Theorem 5.10.
This implies
Theorem 5.15.
Let $\mathcal{O}$ be an order of a number field $K$. Then it can be effectively decided whether there is $\alpha$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$.
Moreover, there are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of $\alpha\in K$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$, and a full system of representatives of those can be effectively determined.
Idea of proof.
Suppose $K$ is effectively given in the form $\mathbb{Q}(\gamma)$, with an algebraic number $\gamma$ of degree $n$. Thus, each element of $K$ has a represention as a $\mathbb{Q}$-linear combinations of $1,\gamma,\ldots,\gamma^{n-1}$, and we can express all computations on $K$ in terms of such representations.
Let the order $\mathcal{O}$ be given by a $\mathbb{Z}$-module basis $1,\theta_{2},\ldots,\theta_{n}$ (with representions as described above). Using Theorem 5.10, one can effectively determine a full system of representatives for the $GL_{2}(\mathbb{Z})$-equivalence classes of those $\alpha\in K$ with $D(\alpha)=D(\mathcal{O})$. To check whether such a representative $\alpha$ satisfies $\mathbb{Z}_{\alpha}=\mathcal{O}$, one can proceed as follows. Verify that $\mathcal{O}\subseteq\mathbb{Z}_{\alpha}$ by checking $\theta_{i}\mathcal{M}_{\alpha}\subseteq\mathcal{M}_{\alpha}$ for $i=2,\ldots,n$. If so, we have in fact $\mathcal{O}=\mathbb{Z}_{\alpha}$ by (5.12). ∎
The rationally monogenic orders introduced above are in fact special cases of invariant orders or invariant rings of binary forms, for which there is now a vast general theory. Although outside the scope of this paper, we give some background on these rings.
Let $A$ be a commutative ring (with $1$), and $a_{0},\ldots,a_{n}\in A$. Then the invariant ring (order if $A=\mathbb{Z}$) associated with $(a_{0},\ldots,a_{n})$, or rather with the binary form $F(X,Y)=a_{0}X^{n}+\cdots+a_{n}Y^{n}$ (but we allow here that $a_{0}=0$ or even $a_{0}=\cdots=a_{n}=0$) is given by the $A$-algebra $A_{F}$ with $A$-module basis $1,\omega_{2},\ldots,\omega_{n}$ satisfying the multiplication table (5.13). This is in fact a commutative, associative $A$-algebra. The name ‘invariant ring’ (invariant order if $A=\mathbb{Z}$) comes from the following invariance property: if $F,G$ are two $GL_{2}(A)$-equivalent binary forms, i.e., $G(X,Y)=uF(aX+bY,cX+dY)$ for some $u\in A^{*}$, $\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)\in GL_{2}(A)$, then $A_{G}\cong A_{F}$ as $A$-algebras. Any ring that is the invariant ring of a binary form is called a binary ring (or binary order if $A=\mathbb{Z}$).
Thus, $\mathbb{Z}_{\alpha}$ is the invariant order of $F_{\alpha}(X,Y):=Y^{n}f_{\alpha}(X/Y)$. In other words, a rationally monogenic order is the invariant order of a primitive, irreducible binary form.
From work of Delone and Faddeev (1940), later extended by Gan, Gross, and Savin (2002) and Deligne (unpublished) (see also section 16.3 of Evertse and Győry (2017)) it follows that for every commutative ring $A$, the map $F\mapsto A_{F}$ gives a one-to-one correspondence between $GL_{2}(A)$-equivalence classes of binary cubic forms in $A[X,Y]$ and isomorphism classes of free cubic $A$-algebras, i.e., commutative, associative, unital $A$-algebras that as an $A$-module are free of rank $3$. Wood (2011) gave a geometric interpretation of invariant rings of binary forms.
@@ 6. Algorithmic resolution of index form equations, application to (multiply) monogenic number fields
The exponential bound (5.10) for the solutions of (5.9) is too large for practical use. In the 1990’s, there were new breakthroughs, leading to the complete resolution of certain index form equations. In fact, practical methods were elaborated for solving equation (5.9) when $|D_{K}|$ is not too large, and the degree $n$ of $K$ is $\leq 6$. Further, (5.9) was solved for some special higher degree number fields $K$ up to about degree $15$ and for some relative extensions of degree $\leq 4$.
@@
As will be seen, in the case $n=3$ equation (5.9) can be reduced to a cubic Thue equation while, in the case $n=4$, to a cubic and some quartic Thue equations, that is to equations of the form
(6.1) $F(x,y)=m\text{ in }x,y\in\mathbb{Z},$
where $m$ is a non-zero integer and $F\in\mathbb{Z}[X,Y]$ is a binary form of degree $3$ or $4$ with pairwise non-proportional linear factors over $\overline{\mathbb{Q}}$. By a general theorem of Thue (1909), every equation of the type (6.1) with degree $\geq 3$ has only finitely many solutions, and Baker (1968b) gave an explicit upper bound for their solutions in terms of $|m|$ and the height and degree of $F$. The best known bound is due to Bugeaud and Győry (1996b). However, in concrete cases this bound is too large for practical use. For solving concrete Thue equations, general practical methods were developed in Pethő and Schulenberg (1987) for $m=1$, and in Tzanakis and de Weger (1989) for arbitrary $m$. Later, these methods were made even more efficient in Bilu and Hanrot (1996, 1999) and Hanrot (1997). Their algorithms are based on Baker’s method and certain reduction and enumeration techniques. Hence we possess efficient algorithms for solving equation (5.9) for $n=3$ and $4$. However, this approach cannot be applied in general to index form equations in number fields $K$ of degree $n>4$, except for $n=6,8,9$ when $K$ has a quadratic or cubic subfield; then equation (5.9) leads to relative cubic or quartic Thue equations.
For $\mathbf{n=3}$, Gaál and Schulte (1989) reduced equation (5.9) to a cubic Thue equation with $m=1$. Then, using the algorithm elaborated for solving cubic Thue equations, they determined all power integral bases of cubic fields $K$ with discriminant $300\leq D_{K}\leq 3137$. Their computations were later extended in Schulte (1989, 1991).
For $\mathbf{n=4}$, Gaál, Pethő and Pohst (1993, 1996) first reduced the equation (5.9) to a cubic Thue equation and a pair of ternary quadratic equations. Then the quadratic equations were themselves reduced to quartic Thue equations. Finally, by means of efficient algorithms for solving such Thue equations, they computed the solutions of equation (5.9) for quartic number fields with not too large discriminant. They obtained several interesting tables on the distribution of minimal indices and about the average behaviour of minimal indices.
6.2. The cases $\mathbf{n=5}$ and $\mathbf{6}$. Refined version of the general approach via unit equations, combined with reduction and enumeration algorithms
@@
For $\mathbf{n\geq 5}$, the approach via Thue equations does not work, in general. For $\mathbf{n=5}$ and $\mathbf{6}$ a refined version of the general approach involving unit equations is needed. Since by (5.7), (5.5) and (5.2) we have for $\alpha\in\mathcal{O}_{K}$ with $K=\mathbb{Q}(\alpha)$,
[…]
where $f_{\alpha}\in\mathbb{Z}[X]$ is the minimal polynomial of $\alpha$, in case of concrete equations (5.9) a refinement of the proof of Theorem 4.2 for irreducible $f_{\alpha}$’s must be combined with some reduction and enumeration algorithms.
The refined version of the general method for solving index form equations (5.9) consists of the following steps:
  • 1.
    Reduction to unit equations but in considerably smaller subfields of the normal closure $G$ of $K$, of which the unit rank is much smaller than that of $G$, i.e., at most n(n1)/21n(n-1)/2-1 (note that the unit rank of $G$ may be as large as n!1n!-1); cf. Győry (1998, 2000). Then in the unit equation corresponding to (4.10), one can write $\varepsilon_{ijk}=\zeta_{ijk}\rho_{1}^{a_{ijk,1}}\cdots\rho_{r}^{a_{ijk,r}}$, with a root of unity $\zeta_{ijk}$ and a fundamental system of units $\rho_{1},\ldots,\rho_{r}$ of bounded height, and in concrete cases one can bound the exponents $|a_{ijk,l}|$ by Baker’s method. Here the estimate of Baker and Wüstholz (1993) for linear forms in logarithms of algebraic numbers is very practical to apply in calculations.
  • 2.
    The bounds in concrete cases are still too large. Hence a reduction algorithm is needed, reducing the Baker’s bound for $|a_{ijk,l}|$ in several steps if necessary by a refined version of the $L^{3}$-algorithm; cf. de Weger (1989), Tzanakis and de Weger (1989), Wildanger (1997) and Gaál and Pohst (1996).
  • 3.
@@
Combining the refined version of the general approach with reduction and enumeration algorithms, for $n=5,6$ and for not too large $|D_{K}|$, Gaál and Győry (1999), resp. Bilu, Gaál and Győry (2004) gave algorithms for determining all power integral bases and hence checking the monogenicity and determining the multiplicity of the monogenicity of $K$.
@@
In the examples below, the authors resolved concrete index form equations of the form (5.9) for $n=3,4,5,6$. The number fields $K$ of degree $n$ are given by irreducible monic polynomials $f(X)\in\mathbb{Z}[X]$, a zero of which generates the corresponding $K$ over $\mathbb{Q}$. In each case all power integral bases in $K$, and therefore the multiplicity of the monogenicity of $K$, denoted by $mm(K)$, are computed by the method outlined above. For the lists of the power integral bases, we refer to the original papers and to Evertse and Győry (2017) and Gaál (2019).
@@
  • $\mathbf{n=6}$, $f(X)=X^{6}-5X^{5}+2X^{4}+18X^{3}-11X^{2}-19X+1$, $mm(K)=45$ (Bilu, Gaál, and Győry, 2004).
We note that from the point of view of computation, the above examples belong to the most difficult cases for $n=3,4,5,$ and $6$, $K$ being in each case totally real with Galois group $S_{n}$. In these cases the number of exponents in the unit equations involved is the largest possible.
Remark. The general procedure outlined above to solve any concrete equation (5.9) for $n=6$ requires considerable CPU-time. In certain special cases (e.g., if $n=6$ and $K$ has a quadratic subfield), there are faster algorithms, see Gaál (2024, 2025). However, some of these algorithms determine only the “small” solutions, and do not exclude the existence of “large" solutions.
For $n\geq 7$, the above mentioned algorithms do not work in general. Then the number of fundamental units, $\rho_{1},\ldots,\rho_{r}$ involved can be $\geq\frac{7\cdot 6}{2}-1=20$ which is too large to use the enumeration algorithm.
@@
For $\mathbf{n=7,8,\ldots}$, give a practical algorithm for solving equation (5.9) in case of any number field $K$ of degree $\mathbf{n}$ with not too large discriminant.